### v3

# The Error in Rayleigh’s Approximative Period

Abstract
We obtain a rigorous a priori upper and lower bounds to the exact period of the celebrated Rayleigh stretched-string differential equation.
1 Introduction
In his monumental treatise The Theory of Sound (see pages 45-46 [2]) Lord Rayleigh proposes the following problem:
A horizontal elastic wire having length $2L_{0}$ and spring constant $\sigma$ is stretched to length $2L$. A body, having a mass $m$ much greater than the mass of the wire, is tied to the center of the wire and is put in motion on a vertical line through the center of the wire. The only forces on the body are produced by tension in the wire (there is no gravity, no damping). It is required to study the motion of the body.
(We have used Agnew’s formulation [1], p. 17-18.) Let $y(t)\equiv y$ be the vertical displacement of the mass $m$ in time $t$, where $y$ is positive, 0, or negative according as the body is above, or at, or below the wire in equilibrium or neutral position.
By the law of Hooke and the definition of a spring constant, $\sigma$, the magnitude of the elastic force, or the tension, $T(y)$, exerted by each half of the stretched string is equal to:
$T(y)$ $=$ $\text{initial tension}+\text{new tension}$
$=$ $\sigma\cdot(L-L_{0})+\sigma\cdot(\sqrt{L^{2}+y^{2}}-L)$
$=$ $\sigma\cdot(\sqrt{L^{2}+y^{2}}-L_{0})$
and therefore the vertical component of the total force, exerted by both halves, which is the only component acting to move the mass $m$ is given by
$\rm{vertical\ component\ of\ T(y)}=-2\sigma\cdot(\sqrt{L^{2}+y^{2}}-L_{0})\frac{y}{\sqrt{L^{2}+y^{2}}},$ (1.1)
where we have employed our sign conventions.
But, by Newton’s second law of motion, that same force, with opposite sign, is given by $m\frac{d^{2}y}{dt^{2}}$.
Therefore, equating the two forces, and bringing the expression for the elastic force on the same side as the force given by Newton’s second law, we find that the algebraic sum of the forces is zero and dividing by $m$, we conclude that the equation of motion of the mass $m$ is given by:
$\boxed{\frac{d^{2}y}{dt^{2}}+\Bigl(\frac{2\sigma}{m}\frac{\sqrt{L^{2}+y^{2}}-L_{0}}{\sqrt{L^{2}+y^{2}}}\Bigr)y=0}$ (1.2)
This differential equation (1.2) is nonlinear and very complicated, and we have never seen an exact treatment of it in the literature.
Rayleigh, himself, simplifies the equation by the following physical reasoning;
“The tension of the string in the position of equilibrium depends on the amount of the stretching to which it has been subjected. In any other position the tension is greater; but we limit ourselves to the case of vibrations so small that the additional stretching is a negligible fraction of the whole. On this condition, the tension may be treated as a constant.
(Our italics.)We just saw that the magnitude of the tension, $T(y)$, (or the force) from each half of the string on $m$ is given by
$T(y)=\sigma(\sqrt{L^{2}+y^{2}}-L_{0})$ (1.3)
provided the stretching is well within elastic limits. So, Rayleigh effectively takes $y=0$ in (1.3) so that the equation (1.2) becomes
$\frac{d^{2}y}{dt^{2}}+\frac{2\sigma}{m}\frac{L-L_{0}}{L}y=0,$ (1.4)
or, using the notation (1.3) with $y=0$, i.e.,
$T:=T(0)=\sigma(L-L_{0})$
we obtain Rayleigh’s approximative differential equation:
$\boxed{\ddot{y}+\frac{2T}{mL}y=0}$ (1.5)
where the “dots” refer to time derivatives. If we assume that the initial displacement at time $t=0$ is $y_{0}$, then the solution to (1.5) is
$\boxed{y(t)=y_{0}\cos\left(\sqrt{\frac{2T}{mL}}t\right)}$ (1.6)
which is simple harmonic motion with period $\overline{P}$ given by:
$\boxed{\overline{P}=\frac{2\pi}{\sqrt{\frac{2T}{mL}}}}$ (1.7)
Rayleigh does not discuss how close his approximate period $\overline{P}$ (1.7) is to the exact period $P$. Indeed, we have been unable to to find any error analysis of Rayleigh’s solution in the literature. Yet, such a famous and classical differential equation merits an investigation into the accuracy of its approximate solution.
In modern mathematics, any approximation, to be acceptable, requires an upper bound on the error to show how good it is, and a lower bound to show how bad it is.
Therefore we offer the following theorem to fill this gap.
Theorem 1.
Let $P$ be the true period of oscilation of the mass $m$. Then Rayleigh’s approximate period $\overline{P}$ overestimates the true period, $P$. Indeed, the following inequalities are valid: If we define the relative error $R$ in the approximation $P\approx\overline{P}$ by
$R:=\frac{P-\overline{P}}{P}$
then
$R<0$
and, in fact
$\boxed{\sqrt{\frac{\frac{1}{L_{0}}-\frac{2}{L+\sqrt{L^{2}+y_{0}^{2}}}}{\frac{1}{L_{0}}-\frac{1}{L}}}<\frac{\overline{P}}{P}<\sqrt{\frac{\frac{1}{L_{0}}-\frac{1}{\sqrt{L^{2}+y_{0}^{2}}}}{\frac{1}{L_{0}}-\frac{1}{L}}}}$ (1.8)
These inequalities show the surprising result that as long as Hooke’s law is applicable, the bound on $R$ only depends on the ratios $\frac{L_{0}}{L}$ and $\frac{y_{0}}{L}$ and not on the mass of the bob nor and Young’s modulus of the material.
We believe the theorem and inequalities to be new.
2 The Formula for the Period
The differential equation (1.2) does not have the variable $t$ nor the first derivative, $\dot{y}$ ,appearing explicity. Therefore, we can transform it into a separable equation for the velocity,
$v:=\dot{y}$ (2.1)
since
$\ddot{y}=\frac{dv}{dt}=\frac{dv}{dy}\frac{dy}{dt}=v\frac{dv}{dy}$ (2.2)
Substituting the right-hand side of (2.2) into (1.2), transposing the expression for the tension to the right-hand side, separating the variables and forming the indefinite integral of both sides we obtain
$\frac{v^{2}}{2}=-\frac{\sigma}{2m}\left(\frac{y^{2}}{2}-L_{0}\sqrt{L^{2}+y^{2}}\right)+C$ (2.3)
where $C$ is the constant of integration. At time $t=0$, the conditions are $y(0)=y_{0}$ and $v(0)=0$. Therefore, evaluating the constant $C$, recalling that $v=\frac{dy}{dt}$, taking the square root of both sides, and using a negative sign because the displacement, $y(t)$, is a decreasing function during the first quarter oscilation, we obtain:
$\frac{dy}{dt}=-\sqrt{\frac{2\sigma L_{0}}{m}(y_{0}^{2}-y^{2})\left(\frac{1}{L_{0}}-\frac{2}{\sqrt{L^{2}+y^{2}}+\sqrt{L^{2}+y_{0}^{2}}}\right).}$ (2.4)
Multiplying the quarter-period by $4$, we obtain the following formula for the full period, $P$, of oscilation of the mass:
Theorem 2.
The true period, $P$, of oscillation of the mass $m$ is given by:
$\boxed{P=4\sqrt{\frac{m}{2\sigma L_{0}}}\int_{0}^{y_{0}}\frac{1}{\sqrt{y_{0}^{2}-y^{2}}\sqrt{\frac{1}{L_{0}}-\frac{2}{\sqrt{L^{2}+y^{2}}+\sqrt{L^{2}+y_{0}^{2}}}}}~dy.}$ (2.5)
This integral is a very complicated elliptic integral and does not permit an elementary exact computation.
Instead, we turn our attention to the upper and lower bounds for the true period, $P$.
3 Upper and Lower Bounds for the True Period
First, we find upper and lower bounds for the period $P$:
Proposition 1.
The true period $P$ is subject to the following inequalities:
$\boxed{\frac{2\pi\sqrt{\frac{m}{2\sigma L_{0}}}}{\sqrt{\frac{1}{L_{0}}-\frac{1}{\sqrt{L^{2}+y_{0}^{2}}}}}\leqslant P\leqslant\frac{2\pi\sqrt{\frac{m}{2\sigma L_{0}}}}{\sqrt{\frac{1}{L_{0}}-\frac{2}{L+\sqrt{L^{2}+y_{0}^{2}}}}}}$ (3.1)
Proof of the bounds on the true period.
The formula for the true period $P$ is ((2.5)):
$P=4\sqrt{\frac{m}{2\sigma L_{0}}}\int_{0}^{y_{0}}\frac{1}{\sqrt{y_{0}^{2}-y^{2}}\sqrt{\frac{1}{L_{0}}-\frac{2}{\sqrt{L^{2}+y^{2}}+\sqrt{L^{2}+y_{0}^{2}}}}}~dy.$
We will find bounds on the second radical in the denominator of the integrand which do not depend on $y$. To this end we note that since $0\leqslant y\leqslant y_{0}$ and $0\leqslant L_{0}\leqslant L$, we conclude that
$L+\sqrt{L^{2}+y_{0}^{2}}\leqslant\sqrt{L^{2}+y^{2}}+\sqrt{L^{2}+y_{0}^{2}}\leqslant 2\sqrt{L^{2}+y_{0}^{2}}$
and after some algebraic manipulation we obtain
$\sqrt{\frac{1}{L_{0}}-\frac{2}{L+\sqrt{L^{2}+y_{0}^{2}}}}\leqslant\frac{1}{\frac{1}{\sqrt{\frac{1}{L_{0}}-\frac{2}{\sqrt{L^{2}+y^{2}}+\sqrt{L^{2}+y_{0}^{2}}}}}}\leqslant\sqrt{\frac{1}{L_{0}}-\frac{1}{\sqrt{L^{2}+y_{0}^{2}}}}$ (3.2)
If we apply this inequality to the denominator in (2.5) and use the results
$\int_{0}^{y_{0}}\frac{dy}{\sqrt{y_{0}^{2}-y^{2}}}=\frac{\pi}{2}$
we obtain the two bounds in the theorem. ∎
Now we use these bounds to obtain the announced bounds on the relative error.
Proof.
Since the relative error
$R:=\frac{P-\overline{P}}{P}\equiv 1-\frac{\overline{P}}{P},$
it is sufficient to bound the fraction $\frac{\overline{P}}{P}.$
First we reduce the formula for Rayleigh’s approximative period:
$\overline{P}=\frac{2\pi}{\sqrt{\frac{2T}{mL}}}=2\pi\sqrt{\frac{mL}{2T}}=\frac{2\pi}{\sqrt{\frac{mL}{2\sigma(L-L_{0})}}}=2\pi\sqrt{\frac{m}{2\sigma}}\sqrt{\frac{L}{L-L_{0}}}$
and therefore, recalling the exact formula for the period $P$, namely (2.5), we obtain
$\frac{\overline{P}}{P}=\frac{2\pi\sqrt{\frac{m}{2\sigma}}\sqrt{\frac{L}{L-L_{0}}}}{4\sqrt{\frac{m}{2\sigma L_{0}}}\int_{0}^{y_{0}}\frac{dy}{\sqrt{y_{0}^{2}-y^{2}}\sqrt{\frac{1}{L_{0}}-\frac{2}{\sqrt{L^{2}+y^{2}}+\sqrt{L^{2}+y_{0}^{2}}}}}.}$
which collapses down to
$\frac{\overline{P}}{P}=\frac{\frac{\pi}{2}\sqrt{\frac{L_{0}L}{L-L_{0}}}}{\int_{0}^{y_{0}}\frac{dy}{\sqrt{y_{0}^{2}-y^{2}}\sqrt{\frac{1}{L_{0}}-\frac{2}{\sqrt{L^{2}+y^{2}}+\sqrt{L^{2}+y_{0}^{2}}}}}.}$ (3.3)
Thus, we must find upper and lower bounds for the denominator in (3.3). But if we combine the bounds in (3.1) with
$\frac{LL_{0}}{L-L_{0}}=\frac{1}{\frac{1}{L_{0}}-\frac{1}{L}}$
and
$\int_{0}^{y_{0}}\frac{dy}{\sqrt{y_{0}^{2}-y^{2}}}=\frac{\pi}{2}$
and cancel the $\frac{\pi}{2}$ in the numerator and the denominator, we obtain the bounds (1.8) in the theorem. ∎
We will finish with some examples.
Example 1.
A wire of length $1152$ milimeters is stretched to a length of $1250$ millimeters. Its spring constant is $1000$ Newtons per meter. A mass of $0.5$ kilograms is attached to its center. The mass is displaced vertically $83$ milimeters and released. It is required to find its period of oscillation.
The exact formula for the period gives $34.1941$ seconds while Rayleigh’s approximative period gives $35.4807$ seconds. Therefore the absolute value of the relative error is $3.4\%$. The bounds in (1.8), after we subtract $1$ give 2.5%⩽3.4%⩽5% which verifies the bounds in the theorem in this case.
The following example explains an anomaly in the applicability of Rayleigh’s approximation.
Example 2.
A steel wire one meter long has a tension of $100$ Newtons. A mass of $0.5$ kilograms is attached to the center and is displaced $2$ centimeters transversely to the wire. It is required to find the period. The spring constant of steel is $157,000$ Newtons per meter.
Rayleigh’s approximate period is $0.2221$ seconds while the true period is $0.18308$ seconds. This gives a relative error of $25\%$!! This example shows that Rayleigh’s approximation can be wildly inaccurate.
Our bounds on the relative error permit us to explain why Rayleigh’s approximation is so bad in this case.
If we use the Taylor expansion of each bound we find
$\frac{1}{2\left(\frac{L}{L_{0}}-1\right)}\left(\frac{y_{0}}{2L}\right)^{2}+O\left(\frac{y_{0}}{2L}\right)^{4}\leqslant|R|\leqslant\frac{1}{\left(\frac{L}{L_{0}}-1\right)}\left(\frac{y_{0}}{2L}\right)^{2}+O\left(\frac{y_{0}}{2L}\right)^{4}$
which shows that to a first approximation, the relative error is directly proportional to the square of the initial fractional displacement and inversely proportional to the initial stretch. Thus, for a given displacement, $\frac{y_{0}}{2L}$, if the stretch is tiny, that is if $L\approx L_{0}$, the relative error will be very large. In this particular case the stretch is $\frac{1}{1570}=0.00064$ meters, and thus the lower bound’s leading term is $15.7\%$ which shows that the true relative error will be even larger. And, indeed, above we found it to be about $25\%$.
Acknowledgment
Support from the Vicerrectoría de Investigación of the University of Costa Rica is acknowledged.